What is relationship between escape and orbital velocity.
1
Expert's answer
2020-10-13T09:51:03-0400

Solution

If "v_e" denotes escape velocity and "v_0" Denote orbital velocity then relationship between escape and orbital velocity can be expressed as below

"\frac{v_e}{v_o}=\sqrt{2}"

At the surface of earth


"v_e=\sqrt{2}v_0" ="1.41v_o"